OpenStudy (anonymous):

she picks 8 people and ask their ages they say:19,27,25,21,44,22,45 and 34 what is the best estimate of the average age of all the people in the tour group round to the nearest whole number

OpenStudy (anonymous):

30/29

OpenStudy (ddos_dragon):

29~

OpenStudy (ddos_dragon):

29.62500 is the exact average,

OpenStudy (anonymous):

Yea

OpenStudy (ddos_dragon):

Round it up to 30.

OpenStudy (anonymous):

so 30?

OpenStudy (anonymous):

I got 29

OpenStudy (ddos_dragon):

Yes, 30.

OpenStudy (anonymous):

When you round up, yes, its 30

OpenStudy (anonymous):

thanks!

OpenStudy (ddos_dragon):

@Stormtrooper You have to round to the nearest whole number 29.62500 rounded is 30 :)

OpenStudy (ddos_dragon):

Remember to medal and close the problem! Have a nice day.

OpenStudy (anonymous):

19 + 27+ 25 + 21 + 44 + 22 + 24 = 203 203/7 = 29.0 But I may be wrong.

OpenStudy (anonymous):

That is supposed to say 22 + 45 not 24

OpenStudy (ddos_dragon):

19+ 27 +25+ 21+ 44+ 22 + 45 + 34= 237

OpenStudy (ddos_dragon):

237/8 = 29.62500

OpenStudy (ddos_dragon):

29.62500 ^ = 30

OpenStudy (anonymous):

Guys, Its 30, Story Over XD